OpenStudy (anonymous):

simplify sec X -sec X sin^2 X

OpenStudy (anonymous):

write out your expand your sin^2 term to (sin)(sin). sec = (sec)(sin)(sin) use the trig identity sec=1/cos on the second sec in your equation sec = (sin)(sin)/cos use the trig identy tan=sin/cos sec=(tan)(sin)